OpenStudy (mikey):

will medal for answer Simplify i12. 1 −1 −i i

OpenStudy (welshfella):

i assume you mean i^12?

OpenStudy (jiteshmeghwal9):

\[i^{4n}=1\]

OpenStudy (mikey):

yes that's what I mean

OpenStudy (jiteshmeghwal9):

\[i^{4*3}=?\]

OpenStudy (welshfella):

i^12 = (i^4)^3

OpenStudy (mikey):

howd he get 1 as the answer

OpenStudy (welshfella):

i^4 = 1 so 1^3 = ?

OpenStudy (welshfella):

- you replace i^4 by 1

OpenStudy (jiteshmeghwal9):

\[i=\sqrt{-1}\]\[i^2=-1\]\[i^{3}=-i\]\[i^4=1\]

OpenStudy (welshfella):

- there you go!

OpenStudy (mikey):

thanks guys

OpenStudy (welshfella):

yw